OpenStudy (joftheworld):

given that (-3,7) is on the graph of f(x) find the corresponding point for the function f(x+5)

OpenStudy (bradely):

f(x+5) =(-3,7+5) =(-3,12) Source: http://www.mathskey.com/question2answer/

OpenStudy (amistre64):

x+5, not y+5

OpenStudy (amistre64):

since f(-3) = 7 f(-3+5) = something that cannot be determined without knowing f(x)